OpenStudy (priyar):

a liquid drop having 6 exçess electrons is kept stationary under electric field of 25.5kV m^-1.The density of liquid is 1.26×10^3 kg m^-3. The radius of the drop is (neglect buoyancy).

OpenStudy (priyar):

@samigupta8 @ganeshie8 @MuH4hA

OpenStudy (priyar):

@agent0smith

OpenStudy (priyar):

@rvc @Michele_Laino @Vincent-Lyon.Fr

OpenStudy (michele_laino):

here on the liquid drop are acting two forces, namely its weight and the electric force, so we can write: \[\Large mg = qE\] where \(q\) is the charge of the liquid drop, and, of course, \(g\) is the gravity. After a substitution, we get: \[\Large \frac{{4\pi }}{3}{R^3}\delta g = 6eE\] where \(\delta\) is the density of the liquid, and \(R\) is the radius of the liquid drop

OpenStudy (samigupta8):

Simply apply qE=mg Q is the charge of 6 electrons

OpenStudy (priyar):

hey I did the exact same thing ..! but answer didn't match.. so i wanted to make sur my method was correct atleast ! thanks @Michele_Laino and @samigupta8 !